OpenStudy (anonymous):

the temperature, T, in degrees celsius, of a yam put into a 200 degrees C oven is given as a function of time, t, in minutes, by: T = a(1-e^kt)+ b. a)if the yam starts at 20 degrees C, find a and b. b)if the temperature of the yam is initially increasing at 2 degrees C per minute, find k.

OpenStudy (anonymous):

k 589002

OpenStudy (anonymous):

duh

OpenStudy (anonymous):

?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

hope i helped

OpenStudy (anonymous):

First off, your equation should be \[T(t)=a(1-e^{-kt})+b\] You need the exponent to be negative, or else the equation blows up at infinity. A) you want to look at the equation at t=0 and t= infinity It's temperature when it starts out, at t=0, is 20ºC, so we can solve for b \[T(0)=20ºC=a(1-\cancel{e^{k*0}}^1)+b \\ \hspace{2.1cm} =a(0)+b\] At t=infinity, the temperature of the cooked yam will be 200ºC \[T(\infty)=200ºC = a(1-\cancel{e^{k*\infty }}^0)+b \] Since we already know b, we can solve for a now. B) So if we know that the yam is increasing it's temperature 2ºC per minute, we know what the temperature of the yam is at t=1 min \[T(1)=22ºC=a(1-e^{-k*1})+b\] And since we know both a and b from the other part, you can solve for k. Remember, to solve exponentials in base e, you take the natural log \[c=e^{-x}\] \[\ln c = \ln e^{-x}\] \[ \ln c = -x\]